OpenStudy (anonymous):

can someone help me diferentiate x^2+4x+3/sqrt of x

OpenStudy (anonymous):

2x+4 -(3/2)x^(-3/2)

OpenStudy (anonymous):

Break it up in parts what is dif of 2x? What is dif of 4?

OpenStudy (anonymous):

I know am suppose to use the quotient rule here but I get stuck @ the algebra part here not the calculus

OpenStudy (laloo):

\[2+3\times(-1\div2)\times(1\div (\sqrt{x})^{3})\]

OpenStudy (anonymous):

How you get that

OpenStudy (laloo):

\[1\div \sqrt{x}=x ^{-1\div2}\] and derivative rule \[x^{n}=n timesx ^{n-1}\] so I got this :)